40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4345     Entry time: Wed Feb 23 16:34:42 2011
Author: valera 
Type: Configuration 
Category:  
Subject: pmc lens staged 

I put the PMC last mode matching lens (one between the steering mirrors) on a translation stage to facilitate the PMC mode matching.

Currently 4% of incident power is reflected by the PMC. But the reflected beam does not look "very professional" on the camera to Rana - meaning there is too much TEM20 (bulls eye) mode in the reflected beam.

I locked the  PMC  on bulls eye mode and measured  the ratio of the TEM20/TEM00 in transmission to be 1.3%. Thus the PMC mode matching is ~99% and the incident beam HOM content is ~3%.

While working on the PMC I found that the source of PMC "blinking" is not the frequency control signal from MC to the laser (the MC servo was turned off) but possibly some oscillation which could be affected even by a small change of the pump current 2.10 A to 2.08 A. I showed this behaviour to Kiwamu and we decided to leave the the current at 2.08 A for now where things look stable and investigate later.

Attachment 1: PMCrefl.JPG  268 kB  Uploaded Wed Feb 23 21:42:23 2011  | Show | Hide all | Show all
Attachment 2: P1070438.JPG  3.006 MB  Uploaded Thu Feb 24 10:49:36 2011  | Hide | Hide all | Show all
P1070438.JPG
Attachment 3: P1070439.JPG  2.896 MB  Uploaded Thu Feb 24 10:54:42 2011  | Hide | Hide all | Show all
P1070439.JPG
ELOG V3.1.3-